OpenStudy (anonymous):

cot(sin^-1 (- sq7/5)

OpenStudy (anonymous):

I'm assuming you wrote \(\cot\left(\sin^{-1}\left(-\dfrac{\sqrt7}{5}\right)\right)\). Suppose we let \(\theta=\sin^{-1}\left(-\dfrac{\sqrt7}{5}\right)\). Then \(\cot \theta\) will be the cotangent of the right triangle with reference angle \(\theta\) that satisfies \(\sin\theta=-\dfrac{\sqrt7}{5}\). Since \(\cot\theta=\dfrac{\cos\theta}{\sin\theta}\), you only need to find the value of \(\cos\theta\).